# Training a Scientific Reasoning Model for Chemistry

### v1

@@ A.2 Group Relative Policy Optimization
The GRPO algorithm is given in Algorithm 1 below. Within it is the following KL divergence estimator [101]:
@@ Appendix C Chemistry RL Dataset Details
C.1 Dataset Provenance
@@
Retrosynthesis tasks used a curated list of experimentally synthesizable molecules. The goal was to propose viable single-step syntheses for these targets. To generate these, we took the fragments from the mcule catalog [^2] https://mcule.com/ and predicted products using the reaction templates from Hartenfeller et al. [102]. Thus, we expected these to be synthesizable. A much larger catalog was used for checking proposed solutions (ZINC20), so that more potential reactions could lead to the products.
@@
C.2 Reward Function Implementation
@@
In tasks that involve submitting a molecule that satisfies constraints, we also do a check on the plausibility of the molecule. See Table 1 for a list of tasks with this check. Aside from assessing if a molecule has valid valence, we check the ring structures and atom fragments. We first take the source molecules for our datasets, which is larger than 640,730 because we did not utilize 100% of ChEMBL or COCONUT. We then applied some filters to ensure the molecules had been synthesized. For example, we required 1 or more assays reported in ChEMBL or a GHS333 Globally Harmonized System of Classification and Labeling of Chemicals categorization being present for molecules from PubChem. The rings from these molecules were isolated using the ring cut method from Pat Walters [103, 104, 105]. The rings were then stored as canonical SMILES in a bloom filter [61]. We then isolated all molecular fragments with radius 2 (2 bonds away) from the molecules and converted them into bit strings similar to ECFP4 fingerprints [86]. These bit strings encode an atom plus its local neighborhood. The bit strings were then stored in a bloom filter. At test time we apply the same ring cuts and fingerprint generation to a proposed molecule. If its rings and fingerprints are all present in the derived bloom filters, we consider the molecule to be reasonable. Otherwise, it is not a reasonable molecule. We use bloom filters because they are highly memory-efficient and fast for checking set membership.
This approach is relatively conservative, because it requires the rings and molecular groups to have been present at least once in a molecule reported in our source datasets. We did experiment with hand-constructed rules, machine learning models, and scores like QED [106], and found them susceptible to reward hacks such as inserting peroxides to satisfy oxygen counts, or hydrazines to increase solubility. We found this check to be essential to ensure plausible molecules are generated. This check is applied at evaluation time as well, and is responsible for rejecting many answers when training the molecule completion and molecular formula tasks.
@@ D.1 Reasoning Quality Filtering
We observed the emergence of reasoning containing typos (made up chemicals), non-english characters (use of languages such as Arabic or Cyrillic), nonsense formatting (blending text with brackets), or ungrounded reasoning (off-the-rails thoughts) as RL progressed. To gauge reasoning quality across training, we employed a LLM judge using the prompt in Section B.3. The judge evaluates reasoning as GREAT, ALRIGHT, or BAD. In practice we found the judgments made by OpenAI GPT 4.1 and Google Gemini 2.5 Pro were interchangeable, and used GPT 4.1 for more favorable rate limits444 Note LLM judges are not 100% reliable, as we observed stray cases where reasoning with non-english characters or typos were labeled as ALRIGHT. Using a regular expression we measured this mistake only occurs in ¡0.1% of judged reasoning traces, so we these results can be trusted as directionally accurate. .
@@
In general, our methodology leaves reasoning unconstrained beyond basic formatting, so it’s intriguing that as task accuracy increases across RL, reasoning flaws begin to appear. Section 5.3 speculates that cognitive behaviors shift in tasks more compatible with reasoning, perhaps there is a similar correlation here, where tasks displaying more reasoning degradation are proportionally less amenable to reasoning models.
@@ D.2 Molecule Quality
[…]
@@ E.1 Task-Specific RL
  • Linear LR warm-up steps: $20$
@@
Since these experiments are relatively short and stable, we did not reset the reference policy during training, but did resume three task-specific runs from a checkpoint (which entails a reference policy reset) to push the model further. Run-specific hyperparameters are detailed in Table S1. DeepSpeed ZeRO Stage 3 [107] was used to shard the model across Nvidia H100 GPUs.
@@ E.2 All-Task RL
The following hyperparameters were used for the final RL phase:
@@
  • Linear LR warm-up steps: $20$
@@ F.2 Cognitive Behavior Counts and Failure Mode Distributions Across Tasks
During evaluation steps performed throughout training, we prompt Llama-3.3-70B-Instruct [108] to analyze each sample generated by our model. For each behavior, we design a custom prompt, following a strategy similar to [25]. Each prompt provides Llama with examples of the target behavior and instructs it to analyze the sample and return the count in a specific format (i.e., <count> [1/2/…] </count>). This procedure enables automatic extraction of behavior counts per sample.
@@ F.3 Advantage-Based Curriculum Ablation
Section 4.3 motivates an advantage-based curriculum; here, we empirically justify its use. In Figure S5, we compare the first few epochs of the reaction prediction specialist (trained with a curriculum) to an identical training run without a curriculum.
@@
Figure S5: RL training dynamics of reaction prediction specialist models, one with and one without an online curriculum. A) the fraction of non-trivial groups seen during training (faint lines are raw data; solid are a 30-step moving average). B) the evaluation set reward, computed every 64 RL steps.
@@ F.4 In-Context Learning
In-context learning (ICL) [2] involves adding examples directly to the prompt at inference time. Also called few-shot, ICL has been shown to improve performance in a range of applications, from property prediction [109, 110, 111] to molecule generation [112, 113, 114]. To build this experiment, we select multiple-choice questions from our dataset and use one of the incorrect choices as context.
@@
F.5 Human evaluation
@@
For the second set of evaluations, we recruited five expert evaluators: three with PhDs in organic chemistry, one with a PhD in chemical engineering, and one PhD candidate in organic chemistry. The evaluators were provided with a rubric to assess the reasoning traces generated by ether0 (see Figure S6a). Each evaluator was given 15 reasoning traces and was compensated \$10 per trace evaluation.
Figure S6: ether0 reasoning trace evaluations by human experts.

### v2

@@ A.2 Group Relative Policy Optimization
The GRPO algorithm is given in Algorithm 1 below. Within it is the following KL divergence estimator [2020_Schulman]:
@@ Appendix C Chemistry RL Dataset Details
C.1 Multiple Choice Question (MCQ) task descriptions
Safety: Select the molecule whose structure most strongly aligns with or deviates from a specified safety-related property, such as toxicity, flammability, or hazard classification.
Scent: Identify the molecule most likely to exhibit a specific olfactory attribute (e.g., meaty, spicy, oily).
Blood-brain barrier: Determine which molecule is or isn’t structurally likely to penetrate the blood-brain barrier based on reference behavior.
Receptor Binding: Identify the molecule whose structure most likely lacks binding affinity or activity for a specified biological receptor target.
ADME: Choose the molecule expected to improve or match a specified absorption, distribution, metabolism, or excretion property based on structural modifications.
Aqueous Solubility: Select the molecule whose structure leads to a targeted increase or decrease in water solubility.
LD50: Select the molecule whose structure corresponds to a specified rat oral LD50 value, reflecting its relative acute toxicity level.
pKa: Select the molecule whose structure aligns with a specified increase, decrease, or target value of pKaH1.
Photoswitches: Select the molecule whose E isomer exhibits a target $\pi$$\pi^{*}$ transition wavelength..
C.2 Dataset Provenance
@@
Retrosynthesis tasks used a curated list of experimentally synthesizable molecules. The goal was to propose viable single-step syntheses for these targets. To generate these, we took the fragments from the mcule catalog [^2] https://mcule.com/ and predicted products using the reaction templates from Hartenfeller et al. [hartenfeller2011collection]. Thus, we expected these to be synthesizable. A much larger catalog was used for checking proposed solutions (ZINC20), so that more potential reactions could lead to the products.
@@
C.3 Reward Function Implementation
@@
In tasks that involve submitting a molecule that satisfies constraints, we also do a check on the plausibility of the molecule. See Table 1 for a list of tasks with this check. Aside from assessing if a molecule has valid valence, we check the ring structures and atom fragments. We first take the source molecules for our datasets, which is larger than 640,730 because we did not utilize 100% of ChEMBL or COCONUT. We then applied some filters to ensure the molecules had been synthesized. For example, we required 1 or more assays reported in ChEMBL or a GHS [^3] Globally Harmonized System of Classification and Labeling of Chemicals categorization being present for molecules from PubChem. The rings from these molecules were isolated using the ring cut method from Pat Walters [walters2022rings, ertl2006quest, usefulrdkitutils]. The rings were then stored as canonical SMILES in a bloom filter [medina2023bloom]. We then isolated all molecular fragments with radius 2 (2 bonds away) from the molecules and converted them into bit strings similar to ECFP4 fingerprints [rogers2010extended]. These bit strings encode an atom plus its local neighborhood. The bit strings were then stored in a bloom filter. At test time we apply the same ring cuts and fingerprint generation to a proposed molecule. If its rings and fingerprints are all present in the derived bloom filters, we consider the molecule to be reasonable. Otherwise, it is not a reasonable molecule. We use bloom filters because they are highly memory-efficient and fast for checking set membership.
This approach is relatively conservative, because it requires the rings and molecular groups to have been present at least once in a molecule reported in our source datasets. We did experiment with hand-constructed rules, machine learning models, and scores like QED [bickerton2012quantifying], and found them susceptible to reward hacks such as inserting peroxides to satisfy oxygen counts, or hydrazines to increase solubility. We found this check to be essential to ensure plausible molecules are generated. This check is applied at evaluation time as well, and is responsible for rejecting many answers when training the molecule completion and molecular formula tasks.
@@ D.1 Reasoning Quality Filtering
We observed the emergence of reasoning containing typos (made up chemicals), non-English characters (use of languages such as Arabic or Cyrillic), nonsense formatting (blending text with brackets), or ungrounded reasoning (off-the-rails thoughts) as RL progressed. To gauge reasoning quality across training, we employed a LLM judge using the prompt in Section B.3. The judge evaluates reasoning as GREAT, ALRIGHT, or BAD. In practice we found the judgments made by OpenAI GPT 4.1 and Google Gemini 2.5 Pro were interchangeable, and used GPT 4.1 for more favorable rate limits [^4] Note LLM judges are not 100% reliable, as we observed stray cases where reasoning with non-English characters or typos were labeled as ALRIGHT. Using a regular expression we measured this mistake only occurs in ¡0.1% of judged reasoning traces, so we these results can be trusted as directionally accurate. .
@@
In general, our methodology leaves reasoning unconstrained beyond basic formatting, so it’s intriguing that as task accuracy increases across RL, reasoning flaws begin to appear.
@@ D.2 Molecule Quality
During the last few steps of GRPO, we further assign a molecule quality bonus reward of 1 to correct answers that also do not contain the above motifs. This is applied to all tasks in Section 2 marked with $\dagger$}.
@@ E.1 Task-Specific RL
  • Linear LR warm-up steps: 2020
@@
Since these experiments are relatively short and stable, we did not reset the reference policy during training, but did resume three task-specific runs from a checkpoint (which entails a reference policy reset) to push the model further. Run-specific hyperparameters are detailed in Table S1. DeepSpeed ZeRO Stage 3 [2020_Rajbhandari] was used to shard the model across Nvidia H100 GPUs.
@@ E.2 All-Task RL
The following hyperparameters were used for the all-task RL phase:
@@
  • Linear LR warm-up steps: 2020
@@
E.3 Safety Alignment
After generalist RL, the model is capable of aiding a user with unsafe chemistry tasks. To improve the safety of the model, we apply a final RL phase for safety alignment. We design the procedure to be independent of model training. We do this to avoid constraining exploration during learning, and to develop a method that is robust enough to be applied as a final phase to any model in our training pipeline.
We construct a specialized dataset comprising questions that a safe model should not respond to. This includes clearly unsafe prompts (e.g., “Design a molecule that is extremely explosive.”), as well as retrosynthesis and edit solubility questions phrased similarly to our training tasks but involving known chemical weapons (e.g., “Suggest a minimal structural modification to ClAsCl that would decrease its solubility by approximately 1 logS unit.”). As a control, we also include benign, open-ended questions that the model should answer correctly but that contain words also seen in the unsafe prompts (e.g., "Design a compound that has no deadly properties.").
To incorporate the new safety behavior into the model, we generate a curated set of prompt-completion examples that include both reasoning traces and the intended refusal response. We then perform a few more steps of GRPO, with both all chemistry tasks and these safety questions. To each group of responses to a safety question, we add the synthetic completion that reflects the desired behavior and assign a reward of 1 to it. In the GRPO objective (Equation 2), we set the importance sampling denominator $\pi_{\theta_{\mathrm{old}}}=1$, following [2021_Libardi].
The following hyperparameters were used for the safety RL phase:
  • Maximum completion length: $4096$
  • Number of training steps: 120
  • Group size: $4$ (non-safety problems) and $5$ (safety problems)
  • Group batch size: $104$
  • GRPO epochs $\mu$: 11
  • Sampling temperature: $1.0$
  • KL penalty weight $\beta$: $0.005$
  • Learning rate: $1\times 10^{-6}$
  • Linear LR warm-up steps: 2020
  • Reference policy reset period $M$: 256 steps
  • Curriculum buffer: ✗
  • Fraction of LLM-rewritten problems: 75%
Figure S2: Performance of ether0 before and after the safety alignment is applied.
@@ F.2 Cognitive Behavior Counts and Failure Mode Distributions Across Tasks
During evaluation steps performed throughout training, we prompt Llama-3.3-70B-Instruct [grattafiori2024llama] to analyze each sample generated by our model. For each behavior, we design a custom prompt, following a strategy similar to [gandhi2025cognitive]. Each prompt provides Llama with examples of the target behavior and instructs it to analyze the sample and return the count in a specific format (i.e., <count> [1/2/…] </count>). This procedure enables automatic extraction of behavior counts per sample.
@@ F.3 Advantage-Based Curriculum Ablation
Section 4.2.1 motivates an advantage-based curriculum; here, we empirically justify its use. In Figure S6, we compare the first few epochs of the reaction prediction specialist (trained with a curriculum) to an identical training run without a curriculum.
@@
Figure S6: RL training dynamics of reaction prediction specialist models, one with and one without an online curriculum. A) the fraction of non-trivial groups seen during training (faint lines are raw data; solid are a 30-step moving average). B) the evaluation set reward, computed every 64 steps.
@@ F.4 In-Context Learning
In-context learning (ICL) [brown2020language] involves adding examples directly to the prompt at inference time. Also called few-shot, ICL has been shown to improve performance in a range of applications, from property prediction [Liu2024-ly, Fifty2023-jz, Schimunek2025-of] to molecule generation [Li2024-qw, Liu2024-ye, Shi2024-yw]. To build this experiment, we select multiple-choice questions from our dataset and use one of the incorrect choices as context.
@@
F.5 Human expert benchmarks
In Table S2, we report the breakdown of human expert performance on our test set.
Task Accuracy
Molecular formula $0.30_{\,0.13}^{\,0.60}$
Functional group $0.13_{\,0.00}^{\,0.30}$
Reaction prediction $0.41_{\,0.08}^{\,0.72}$
Solubility edit $0.05_{\,0.00}^{\,0.12}$
Retrosynthesis 0.00 0.00 $_{\,0.00}^{\,0.00}$
Safety $0.40_{\,0.28}^{\,0.52}$
ADME $0.32_{\,0.24}^{\,0.48}$
LD50 $0.43_{\,0.24}^{\,0.64}$
pKa $0.35_{\,0.20}^{\,0.48}$
Table S2: Human expert performance on the test set. Four contractors were tasked with solving each question, with no tools besides ChemDraw. We report the average accuracy, as well as the minimum and maximum scores, as sub- and super-scripts respectively.
F.6 Human evaluation
@@
For the second set of evaluations, we recruited another group of expert evaluators: three with PhDs in organic chemistry, and one with a PhD in chemical engineering. The evaluators were provided with a rubric to assess the reasoning traces generated by ether0 and DeepSeek-R1 (see Figure S7). Each evaluator was given 30 reasoning traces from both model (15 from each). Compensation was \$10 per completed trace evaluation. Full comparison results are shown in Table S3. More experts disagreed that DeepSeek-r1 demonstrated non-contrived and faithful reasoning, although they noted it showed more extensive exploration. This is unsurprising, as DeepSeek-r1’s reasoning traces were much longer than ether0’s.
Figure S7: Four expert evaluators were provided with this rubric to assess the “quality” of 15 traces from ether0 and 15 traces from DeepSeek-R1.
Non-Contrived / Genuine S. Agree Agree Neutral Disagree S. Disagree
DeepSeek-r1 28% 23% 7% 8% 33%
ether0 45% 23% 12% 15% 5%
Faithful S. Agree Agree Neutral Disagree S. Disagree
DeepSeek-r1 33% 25% 8% 20% 13%
ether0 50% 23% 10% 13% 3%
Explorative S. Agree Agree Neutral Disagree S. Disagree
DeepSeek-r1 20% 32% 18% 13% 17%
ether0 13% 28% 35% 5% 18%
Table S3: Expert evaluation of reasoning traces generated by ether0 and DeepSeek-R1.